2 Direct material Advertising expense Depreciation on factory building Direct labor: wages Cost of finished goods inventory at year-en Indirect labor: wages Production supervisors sala Service department costs Direct labor: fringe benefits Indirect labor: fringe benefits Fringe benefits for production supervisor Total overtime premiums paid Cost of idle time: production Administrative costs $2,300,800 97,060 116,000 515,000 115,000 2.5 points 42 47,800 100,a00 95,900 31,000 9,000 55,000 40,000 50,000 15,000 5,000 10,000 ees Rental of office space for sales personnel Sales commissions Product promotion cos Al I services are provided to manufacturing department SCost of idle time isa tThe rental of sales space was material not direct-labor wages given above s made necessary when the sales ofices were converted to storage space for fa Required 1. Compute each of the following costs for the year just enc 8 Answer is not complet a Total prime costs otal manufacturing overhe costs c Total conversion cost d. Total product costs e Total period costs 300,000 Nex
Answer #1
a. Total prime costs $       2,910,000
b. Total manufacturing overhead costs $          540,000
c. Total conversion costs $       1,150,000
d. Total product costs $       3,450,000
e. Total period costs $          277,000

Working:

DM DL MOH Period cost
Direct material 2300000
Advertising expense 97000
Depreciation on factory building 116000
Direct labor: wages 515000
Indirect labor: wages 142000
Production supervisor's salary 47000
Service department costs 100000
Direct labor: fringe benefits 95000
Indirect labor: fringe benefits 31000
Fringe benefits for production supervisor 9000
Total overtime premiums paid 55000
Cost of idle time: production employees 40000
Administrative costs 150000
Rental of office space for sales personnel 15000
Sales commissions 5000
Product promotion costs 10000
Total costs $ 2300000 610000 540000 277000

Prime cost = DM + DL = 2300000 + 610000 = $2910000

Conversion cost = DL + MOH = 610000 + 540000 = $1150000

Product cost = DM + DL + MOH = 2300000 + 610000 + 540000 = $3450000